OpenStudy (anonymous):

HELP!!! Im totally confused! factor this monster! 3x^3+2x^2-5x+9x^2y+6xy-15y HELP!!! Im totally confused! factor this monster! 3x^3+2x^2-5x+9x^2y+6xy-15y @Mathematics

OpenStudy (anonymous):

(x-1)(3x+5)(x+3y)

OpenStudy (anonymous):

and how did u get that?

OpenStudy (anonymous):

common knowledge don't know how to explain

OpenStudy (anonymous):

yea...i kinda needed to knw how to do it...

OpenStudy (anonymous):

\((3x^3+2x^2-5x)+(9x^2y+6xy-15y) \implies x(3x^2+2x-5)+3y(3x^2+2x-5) \). You czn see here that the parentheses \((3x^2+2x-5)\) is common, so we can take it as a common factor here. It can also be factorized to \((x-1)(3x+5)\) . Our expression will be: \( x(3x^2+2x-5)+3y(3x^2+2x-5)\) \(=(x+3y)(3x^2+2x-5)=(x+3y)(x-1)(3x+5).\)

OpenStudy (anonymous):

oh..so u took out the x in one of them?

OpenStudy (anonymous):

Yeah, I took out the x out of the first parenthesis and 3y out of the second.

OpenStudy (anonymous):

OH...okay..i get it now! thank u!

OpenStudy (anonymous):

common knowledge = wolframalpha http://www.wolframalpha.com/input/?i=factor+3x%5E3%2B2x%5E2-5x%2B9x%5E2y%2B6xy-15y